OpenStudy (anonymous):

A rewitten partial fraction (3x+2)/(x^2-x-12) includes which? 1/x+3 or 1/x-4 or 2/x-4

OpenStudy (anonymous):

\[\frac{3x+2}{x^2-x-12}=\frac{3x+2}{(x-4)(x+3)}=\frac{A}{x-4}+\frac{B}{x+3}\] \[3x+2=(A+B)x+3A-4B\\ \begin{cases}A+B=3\\3A-4B=2\end{cases}\] Solve for A and B.

OpenStudy (anonymous):

so 2/x-4 is not a factor at all?

OpenStudy (anonymous):

That depends. What are you getting for A and B?